### v3

@@ 3.4 Finite Difference Monte Carlo estimators
$\begin{split}\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}=\frac{1}{N}\sum_{n=1}^{N}\mathsf{APURE}_{\boldsymbol{\zeta}^{(n)}}^{\mathcal{E}}\quad\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}=\frac{1}{N}\sum_{n=1}^{N}\mathsf{APURE}_{\boldsymbol{\zeta}^{(n)}}^{\mathcal{P}}\end{split}$ (39)
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathcal{E}(\widehat{\boldsymbol{\mathsf{X}}},\overline{\boldsymbol{\mathsf{X}}})\quad\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathcal{P}(\widehat{\boldsymbol{\mathsf{X}}},\overline{\boldsymbol{\mathsf{X}}}).\end{split}$ (40)
@@ 4 Application to piecewise linear estimation
The purpose of this section is twofold: first, to assess the ability of the robustified Finite Difference Monte Carlo risk estimates deriving from Theorem 2 and Proposition 1, to approximate faithfully the true estimation and prediction risks; second, to demonstrate numerically that the hyperparameters selected by minimizing these risk estimates yields accurate estimates of the reproduction coefficients from observations following the generalized time-varying autoregressive model with data-dependent Poisson noise (11). To that aim, intensive simulations are run on synthetic data generated according to (7) and (11).
@@ 4.1 Synthetic data
To prepare for the application to epidemiological indicator estimation, developed in Section 5, the ground truth reproduction coefficient $\overline{\boldsymbol{\mathsf{X}}}$ is designed piecewise linear [1, 59], imitating temporal evolution of the reproduction number of COVID-19 observed from real-world data [21]. All synthetic data in this section are of length $T=70$ and share the same ground truth, represented by the deep blue curve in Figure 1, bottom plot, alternating expansion and recession phases, as represented by the blue areas indicating exponential growth period characterized by $\overline{\mathsf{X}}_{t}>1$. The initial state is set according to the observed real COVID-19 infection counts during the imitated period at $\mathsf{Y}_{0}=3395$. Mimicking the epidemiological model proposed in [17] particularized to COVID-19 pandemic, the memory functions are chosen linear, as described in Example 1, with a constant memory horizon of $\tau=25$ and a sequence $\{\psi_{s}\}_{s=1}^{\tau}$ chosen as the daily discretization of the serial interval distribution of COVID-19 modeled as a Gamma distribution of mean $6.6$ days and standard deviation $3.5$ days [9, 66]. The scale parameter of the data-dependent Poisson noise is constant through time, that is $\forall t,\alpha_{t}=\alpha>0$. Seven values of $\alpha$ logarithmically spaced between $\alpha=10^{2}$, corresponding to low noise level, to a very high noise level of $\alpha=10^{4.5}$, are explored. Figure 2 provides in top row examples of synthetic observations with the same underlying ground truth reproduction coefficient, displayed in blue in the second row plots, for three scale parameter values, corresponding, from first to third columns, to low $\alpha=10^{2}$, medium $\alpha=10^{3}$ and high $\alpha=10^{4}$ noise levels.
@@
Figure 2: Estimation of the reproduction coefficient from synthetic generalized time-varying autoregressive Poisson observations using the variational estimator (41) coupled with the oracle-based hyperparameter selection strategy (4). First row: One realization of synthetic observations drawn following the setup described in Section 4.1. Second row: Underlying ground truth reproduction coefficient in blue, and estimates obtained from the variational estimator (41) combined with the oracle-based regularization parameter selection strategy, for the two ground truth-dependent oracles $\mathcal{P}^{\circ}$ and $\mathcal{E}^{\circ}$ in red and green respectively, and for the two data-driven oracles $\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ and $\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ in pink and orange respectively. Third and fourth rows: Robustified data-driven prediction and estimation risk estimates computed on a logarithmic grid of regularization parameters $\lambda$, accompanied with their Gaussian confidence regions computed from the $N=10$ Monte Carlo vector realizations, and exact prediction and estimation errors; vertical dashed lines indicate optimal hyperparameters selected for each oracle.
@@ 4.2 Estimation strategy
where $\mathcal{D}$ is the Kullback-Leibler divergence customized to the generalized time-varying autoregressive Poisson model as described in Equation (18), $\textbf{D}_{2}$ is the discrete Laplacian, which, combined with the $\ell_{1}$ norms favors piecewise linear behavior of the estimate, and $\lambda>0$ is a regularization parameter balancing the data-fidelity and the penalization [1, 59]. As the minimization problem (41) is convex but nonsmooth, it is solved using the Chambolle-Pock proximal algorithm [10, 1, 59] leveraging the closed-form expression of the proximity operator of the Kullback-Leibler divergence of Equation (18[1].
@@
which, by definition of $\mathcal{E}$ and $\mathcal{P}$, are ground truth-dependent unbiased estimates of the estimation and prediction risk respectively; and the proposed $\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}$ and $\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}$ introduced in Section 3, which are fully data-driven. Oracles are minimized through exhaustive search over a logarithmic grid of $\lambda$ from $10^{-2}\times\mathrm{std}(\boldsymbol{\mathsf{Y}})$ to $10^{4}\times\mathrm{std}(\boldsymbol{\mathsf{Y}})$, providing four hyperparameter choices $\lambda_{\mathcal{O}}$, and resulting in four estimates $\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\lambda_{\mathcal{O}})$.6 Numerical simulations aim at assessing the quality of the estimates obtained from the proposed data-driven oracles compared to estimates based on ground truth oracles, which are not usable in real-world problems, such as epidemic monitoring described in Section 5.
Metric Minimized oracle $\log_{10}\alpha=2$ $\log_{10}\alpha=2.5$ $\log_{10}\alpha=3$ $\log_{10}\alpha=3.5$ $\log_{10}\alpha=4$ $\log_{10}\alpha=4.5$
$\mathsf{MMSE}\pm\mathsf{CI}$ Estimation error $\mathcal{E}^{\circ}$ $0.51\pm 0.02$ $0.57\pm 0.05$ $1.06\pm 0.20$ $1.23\pm 0.40$ $1.76\pm 0.52$ $1.37\pm 0.26$
Prediction error $\mathcal{P}^{\circ}$ $0.53\pm 0.02$ $0.59\pm 0.05$ $1.21\pm 0.32$ $2.38\pm 2.00$ $1.89\pm 0.52$ $1.50\pm 0.27$
$\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ $0.53\pm 0.02$ $0.59\pm 0.05$ $1.20\pm 0.33$ $1.40\pm 0.44$ $2.17\pm 0.68$ $1.53\pm 0.29$
$\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ $0.53\pm 0.02$ $0.59\pm 0.05$ $2.03\pm 1.28$ $4.86\pm 4.10$ $7.42\pm 4.42$ $23.98\pm 17.82$
$\mathsf{Bias}$$\mathsf{Variance}$ Estimation error $\mathcal{E}^{\circ}$ 0.490.490.490.490.020.020.020.02 0.500.500.500.500.060.060.060.06 0.730.730.730.730.330.330.330.33 0.550.550.550.550.680.680.680.68 0.850.850.850.850.910.910.910.91 0.550.550.550.550.820.820.820.82
Prediction error $\mathcal{P}^{\circ}$ 0.490.490.490.490.040.040.040.04 0.500.500.500.500.090.090.090.09 0.660.660.660.660.550.550.550.55 0.820.820.820.821.561.561.561.56 0.770.770.770.771.111.111.111.11 0.480.480.480.481.021.021.021.02
$\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ 0.500.500.500.500.030.030.030.03 0.530.530.530.530.070.070.070.07 0.700.700.700.700.500.500.500.50 0.560.560.560.560.840.840.840.84 0.910.910.910.911.261.261.261.26 0.570.570.570.570.960.960.960.96
$\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ 0.500.500.500.500.030.030.030.03 0.510.510.510.510.080.080.080.08 0.720.720.720.721.311.311.311.31 0.950.950.950.953.913.913.913.91 0.990.990.990.996.436.436.436.43 1.471.471.471.4722.5122.5122.5122.51
@@ 4.3 Performance evaluation
For each scale parameter explored, performances are evaluated on a collection of $Q=20$ independent synthetic observations $\{\boldsymbol{\mathsf{Y}}^{(q)},\,q=1,\ldots,Q\}$ generated following the setup of Section 4.1. For $\mathcal{O}\in\{\mathcal{E}^{\circ},\mathcal{P}^{\circ},\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N},\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}\}$, the hyperparameters selected through the oracle strategy applied to the $q$th realization are denoted [^6] The dependency of $\lambda_{\mathcal{O}}$ in $\boldsymbol{\mathsf{Y}}$ is omitted for readability.
@@
The number of realizations $Q$ over which the performance is computed has been chosen after observing that all aforementioned metrics together with their confidence intervals had stabilized after $Q=20$.
@@ 4.4 Results
Figure 2 compares the four oracle-based estimation strategies on one realization of generalized time-varying autoregressive Poisson synthetic observations for three values of the scale parameter, corresponding to low, medium and high noise levels. Synthetic observations and memory terms are displayed in the first row; the underlying ground truth reproduction coefficient and its estimates obtained using the four different oracles are plotted in blue on the second row; third (resp. fourth) row compares the prediction (resp. estimation) ground truth and data-dependent oracles, and the associated optimal hyperparameters.
@@ 5 Application to epidemiology
A major motivation of the present work lies in the need for data-driven hyperparameter fine-tuning strategies for recently proposed COVID-19 reproduction estimators leveraging the variational framework sketched in Equation (2[1, 59, 21].
@@ 5.1 Weekly scaled Poisson epidemiological model
The considered epidemiological model, briefly introduction in Section 1, in Equation (5), was initially proposed in [17] and states that, conditionally to past infection counts $\mathsf{Z}_{1},\ldots,\mathsf{Z}_{t-1}$, the number of new infections at time $t$, denoted $\mathsf{Z}_{t}$, follows a Poisson distribution of intensity equal to the product of the effective reproduction number at time $t$, $\mathsf{R}_{t}$, and the global infectiousness $\Phi_{t}(\boldsymbol{\mathsf{Z}})=\sum_{s\geq 1}\varphi_{s}\mathsf{Z}_{t-s}$ with $\varphi$ the serial interval distribution. [^8] The serial interval is the random delay between primary and secondary infections; its distribution encodes the typical time scales of the propagation. Although very accurate for a posteriori consolidated data [17], the daily version of this model appeared not suitable for real-time COVID-19 daily infection counts, an example of which is displayed in Figure LABEL:fig:daily, gray curve. Indeed, COVID-19 data are severely corrupted by administrative noise, taking the form of missing counts during week-ends and holidays, large cumulative counts on Mondays, pseudo-seasonalities, erroneous samples, to name but a few. This noise has two major effects: first, reported COVID-19 daily infections counts contain outlier values, notably during week-ends as illustrated in the gray curve in Figure LABEL:fig:daily; second, the variance of the reported counts reflects both the intrinsic variance of the propagation process and the additional variance induced by the fluctuating administrative delays and errors. To circumvent the presence of outlier samples, a classical strategy in epidemiology is to consider aggregated data [25, 13, 53], e.g., at the scale of the week as illustrated in Figure LABEL:fig:daily, black curve, which is far smoother and hence more realistic from an epidemiological point of view. To go further in the modeling of infection counts, the present work proposes to account for the increased variance observed in real COVID-19 infection counts through a scale parameter $\alpha>1$ constant in time. Up to the authors’ knowledge, scaled Poisson distributions have not yet been explored in the epidemiology literature, making this an original contribution. Altogether, the proposed weekly scaled Poisson epidemiological model writes

### v4

@@ 3.4 Finite Difference Monte Carlo estimators
$\begin{split}\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}=\frac{1}{N}\sum_{n=1}^{N}\mathsf{APURE}_{\boldsymbol{\zeta}^{(n)}}^{\mathcal{E}}\quad\text{and}\quad\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}=\frac{1}{N}\sum_{n=1}^{N}\mathsf{APURE}_{\boldsymbol{\zeta}^{(n)}}^{\mathcal{P}}\end{split}$ (38)
@@
$\begin{split}\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathcal{E}(\widehat{\boldsymbol{\mathsf{X}}},\overline{\boldsymbol{\mathsf{X}}})\quad\text{and}\quad\mathbb{E}_{\boldsymbol{\mathsf{Y}},\boldsymbol{\zeta}}\left[\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}\right]\underset{\boldsymbol{\alpha}\rightarrow\boldsymbol{0}}{=}\mathcal{P}(\widehat{\boldsymbol{\mathsf{X}}},\overline{\boldsymbol{\mathsf{X}}}).\end{split}$ (39)
@@ 4 Application to piecewise linear estimation
The purpose of this section is twofold: first, to assess the ability of the robustified Finite Difference Monte Carlo risk estimates derived from Theorem 2 and Proposition 1, to approximate faithfully the true estimation and prediction risks; second, to demonstrate numerically that the hyperparameters selected by minimizing these risk estimates yield accurate estimates of the reproduction coefficients from observations following the generalized time-varying autoregressive model with data-dependent Poisson noise (10). To that aim, intensive simulations are run on synthetic data generated according to (6) and (10).
@@ 4.1 Synthetic data
To prepare for the application to epidemiological indicator estimation, developed in Section 5, the ground truth reproduction coefficient $\overline{\boldsymbol{\mathsf{X}}}$ is designed piecewise linear [1, 59], imitating temporal evolution of the reproduction number of COVID-19 observed from real-world data [21]. All synthetic data in this section are of length $T=300$ and share the same ground truth, represented by the deep blue curve in Figure 1, bottom plot, alternating expansion and recession phases, as represented by the blue areas indicating exponential growth period characterized by $\overline{\mathsf{X}}_{t}>1$. The initial state is set according to the observed real COVID-19 infection counts during the imitated period at $\mathsf{Y}_{0}=3395$. Mimicking the epidemiological model proposed in [17] particularized to COVID-19 pandemic, the memory functions are chosen linear, as described in Example 1, with a constant memory horizon of $\tau=25$ and a sequence $\{\psi_{s}\}_{s=1}^{\tau}$ chosen as the daily discretization of the serial interval distribution of COVID-19 modeled as a Gamma distribution of mean $6.6$ days and standard deviation $3.5$ days [9, 66]. The scale parameter of the data-dependent Poisson noise is constant through time, i.e., $\forall t,\,\alpha_{t}=\alpha>0$. Seven values of $\alpha$ logarithmically spaced between $\alpha=10^{2}$, corresponding to low noise level, to a very high noise level of $\alpha=10^{5}$, are explored. Figure 2 provides in top row examples of synthetic observations with the same underlying ground truth reproduction coefficient, displayed in blue in the second row plots, for three scale parameter values, corresponding, from first to third columns, to low $\alpha=10^{2}$, medium $\alpha=10^{3}$ and high $\alpha=10^{4}$ noise levels.
@@
Figure 2: Estimation of the reproduction coefficient from synthetic generalized time-varying autoregressive Poisson observations using the variational estimator (40) coupled with the oracle-based hyperparameter selection strategy (4). First row: One realization of synthetic observations drawn following the setup described in Section 4.1. Second row: Underlying ground truth reproduction coefficient in blue, Maximum Likelihood estimate (21) in light blue, state-of-the-art estimator EpiEstim (42) derived for reproduction number estimation in epidemiology [17, 74] in dark gray, and estimates obtained from the variational estimator (40) combined with the oracle-based regularization parameter selection strategy, for the two ground truth-dependent oracles $\mathcal{P}^{\circ}$ and $\mathcal{E}^{\circ}$ in red and green respectively, and for the two data-driven oracles $\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ and $\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ in pink and orange respectively. Third and fourth rows: Robustified data-driven prediction and estimation risk estimates computed on a logarithmic grid of regularization parameters $\lambda$, accompanied with their Gaussian confidence regions computed from the $N=10$ Monte Carlo vector realizations, and exact prediction and estimation errors; vertical dashed lines indicate optimal hyperparameters selected for each oracle.
@@ 4.2 Estimation strategy
where $\mathcal{D}$ is the Kullback-Leibler divergence customized to the generalized time-varying autoregressive Poisson model as described in Equation (17), $\textbf{D}_{2}$ is the discrete Laplacian, which, combined with the $\ell_{1}$-norm favors piecewise linear behavior of the estimate, and $\lambda>0$ is a regularization parameter balancing the data-fidelity and the penalization [1, 59]. As the minimization problem (40) is convex but nonsmooth, it is solved using the Chambolle-Pock proximal algorithm [10, 1, 59] leveraging the closed-form expression of the proximity operator of the Kullback-Leibler divergence of Equation (17[1].
@@
which, by definition of $\mathcal{E}$ and $\mathcal{P}$, are ground truth-dependent unbiased estimates of the estimation and prediction risk respectively; and the proposed $\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N}$ and $\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}$ introduced in Section 3, which are fully data-driven. Oracles are minimized through exhaustive search over a logarithmic grid of $\lambda$ from $10^{-2}\times\mathrm{std}(\boldsymbol{\mathsf{Y}})$ to $10^{4}\times\mathrm{std}(\boldsymbol{\mathsf{Y}})$, providing four hyperparameter choices $\lambda_{\mathcal{O}}$, and resulting in four estimates $\widehat{\boldsymbol{\mathsf{X}}}(\boldsymbol{\mathsf{Y}};\lambda_{\mathcal{O}})$. [^6] The dependency of $\lambda_{\mathcal{O}}$ in $\boldsymbol{\mathsf{Y}}$ is omitted for readability. Numerical simulations aim at assessing the quality of the estimates obtained from the proposed data-driven oracles compared to estimates based on ground truth oracles, which are not usable in real-world problems, such as epidemic monitoring described in Section 5.
Furthermore, these four estimates are compared to the Maximum Likelihood estimate of Equation (21) and to EpiEstim [^7] https://github.com/jstockwin/EpiEstimApp [17, 74], the state-of-the-art reproduction number estimator in epidemiology, obtained as the a posteriori mean of a Bayesian model composed of the Poisson autoregressive likehood (17) together with a conjugated gamma prior of shape parameter $a=1$ and scale parameter $b=5$ on the reproduction coefficient and smooth behavior constraint, yielding
$\forall t\in\{1,\ldots,T\},\quad\widehat{\mathsf{X}}_{t}^{\mathrm{EpiEstim}}=\frac{a+\sum_{s=0}^{\delta-1}\mathsf{Y}_{t-s}}{1/b+\sum_{s=0}^{\delta-1}\Psi_{t-s}(\boldsymbol{\mathsf{Y}})}$ (42)
where $\delta=7$ mimics the counts and infectiousness averaging over seven days which enforces temporal regularity [17].
Metric Estimator $\log_{10}\alpha=2$ $\log_{10}\alpha=2.5$ $\log_{10}\alpha=3$ $\log_{10}\alpha=3.5$ $\log_{10}\alpha=4$ $\log_{10}\alpha=4.5$ $\log_{10}\alpha=5$
$\mathsf{MMSE}\pm\mathsf{CI}$ Max. likelihood (21) $1.25\pm 0.08$ $3.17\pm 0.34$ $20.58\pm 9.52$ $28.38\pm 8.87$ $56.25\pm 14.40$ $83.30\pm 32.98$ $228.00\pm 374.80$
EpiEstim [17, 74] $1.26\pm 0.04$ $1.47\pm 0.09$ $3.96\pm 1.33$ $5.80\pm 1.93$ $8.59\pm 1.64$ $11.20\pm 3.82$ $18.83\pm 9.18$
Estimation error $\mathcal{E}^{\circ}$ $0.51\pm 0.02$ $0.57\pm 0.05$ $1.06\pm 0.20$ $1.23\pm 0.40$ $1.76\pm 0.52$ $1.37\pm 0.26$ $2.35\pm 1.51$
Prediction error $\mathcal{P}^{\circ}$ $0.53\pm 0.02$ $0.59\pm 0.05$ $1.21\pm 0.32$ $2.38\pm 2.00$ $1.89\pm 0.52$ $1.50\pm 0.27$ $3.03\pm 1.84$
$\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ $0.53\pm 0.02$ $0.59\pm 0.05$ $1.20\pm 0.33$ $1.40\pm 0.44$ $2.17\pm 0.68$ $1.53\pm 0.29$ $8.80\pm 12.90$
$\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ $0.53\pm 0.02$ $0.59\pm 0.05$ $2.03\pm 1.28$ $4.86\pm 4.10$ $7.42\pm 4.42$ $23.98\pm 17.82$ $42.08\pm 44.06$
$\mathsf{Bias}$$\mathsf{Variance}$ Max. likelihood (21) 0.520.520.730.73 0.590.592.582.58 1.691.6918.8818.88 1.941.9426.4426.44 3.963.9652.2952.29    4.504.5078.8078.80 15.8815.88262.10262.10
EpiEstim [17, 74] 1.151.150.110.11 1.161.160.310.31 1.331.332.632.63 1.361.364.454.45 1.601.606.996.99 1.301.309.909.90 1.481.4817.3517.35
Estimation error $\mathcal{E}^{\circ}$ 0.490.490.020.02 0.500.500.060.06 0.730.730.330.33 0.550.550.680.68 0.850.850.910.91 0.550.550.820.82 0.980.981.371.37
Prediction error $\mathcal{P}^{\circ}$ 0.490.490.040.04 0.500.500.090.09 0.660.660.550.55 0.820.821.561.56 0.770.771.111.11 0.480.481.021.02 0.650.652.382.38
$\overline{\mathsf{APURE}^{\mathcal{P}}_{\boldsymbol{\zeta}}}^{N}$ 0.500.500.030.03 0.530.530.070.07 0.700.700.500.50 0.560.560.840.84 0.910.911.261.26 0.570.570.960.96 1.041.047.757.75
$\overline{\mathsf{APURE}^{\mathcal{E}}_{\boldsymbol{\zeta}}}^{N}$ 0.500.500.030.03 0.510.510.080.08 0.720.721.311.31 0.950.953.913.91 0.990.996.436.43     1.471.4722.5122.51 2.382.3839.7039.70
@@ 4.3 Performance evaluation
For each scale parameter explored, performances are evaluated on a collection of $Q=20$ independent synthetic observations $\{\boldsymbol{\mathsf{Y}}^{(q)},\,q=1,\ldots,Q\}$ generated following the setup of Section 4.1. For $\mathcal{O}\in\{\mathcal{E}^{\circ},\mathcal{P}^{\circ},\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{E}}}^{N},\overline{\mathsf{APURE}_{\boldsymbol{\zeta}}^{\mathcal{P}}}^{N}\}$, the hyperparameters selected through the oracle strategy applied to the $q$th realization are denoted6
@@
The number of realizations $Q$ over which performance are computed has been chosen after observing that all aforementioned metrics together with their confidence intervals had stabilized after $Q=20$ realizations.
@@ 4.4 Results
Figure 2 compares the four proposed oracle-based estimation strategies, the Maximum Likelihood (21) and the state-of-the-art EpiEstim [17, 74] estimators on one realization of generalized time-varying autoregressive Poisson synthetic observations for three values of the scale parameter, corresponding to low, medium and high noise levels. Synthetic observations and memory terms are displayed in the first row; the underlying ground truth reproduction coefficient and its estimates obtained using the four different oracles, Maximum Likelihood strategy and EpiEstim estimator are plotted on the second row; third (resp. fourth) row compares the prediction (resp. estimation) ground truth and data-dependent oracles, and the associated optimal hyperparameters.
@@ 5 Application to epidemiology
A major motivation of the present work is the need for data-driven hyperparameter fine-tuning strategies for recently proposed COVID-19 reproduction estimators leveraging the variational framework sketched in Equation (2[1, 59, 21].
@@ 5.1 Weekly scaled Poisson epidemiological model
The considered epidemiological model, briefly introduction in Section 1, in Equation (5), was initially proposed in [17] and states that, conditionally to past infection counts $\mathsf{Z}_{1},\ldots,\mathsf{Z}_{t-1}$, the number of new infections at time $t$, denoted $\mathsf{Z}_{t}$, follows a Poisson distribution of intensity equal to the product of the effective reproduction number at time $t$, $\mathsf{R}_{t}$, and the global infectiousness $\Phi_{t}(\boldsymbol{\mathsf{Z}})=\sum_{s\geq 1}\varphi_{s}\mathsf{Z}_{t-s}$ with $\varphi$ the serial interval distribution. [^9] The serial interval is the random delay between primary and secondary infections; its distribution encodes typical propagation time scale. Although very accurate for a posteriori consolidated data [17], the daily version of this model appeared not suitable for real-time COVID-19 daily infection counts, an example of which is displayed in Figure LABEL:fig:daily, gray curve. Indeed, COVID-19 data are severely corrupted by administrative noise, taking the form of missing counts during week-ends and holidays, large cumulative counts on Mondays, pseudo-seasonalities, erroneous samples, to name but a few. This noise has two major effects: first, reported COVID-19 daily infections counts contain outlier values, notably during week-ends as illustrated in the gray curve in Figure LABEL:fig:daily; second, the variance of the reported counts reflects both the intrinsic variance of the propagation process and the additional variance induced by the fluctuating administrative delays and errors. To circumvent the presence of outlier samples, a classical strategy in epidemiology is to consider aggregated data [25, 13, 53], e.g., at the scale of the week as illustrated in Figure LABEL:fig:daily, black curve, which is far smoother and hence more realistic from an epidemiological point of view. To go further in the modeling of infection counts, the present work proposes to account for the increased variance observed in real COVID-19 infection counts through a scale parameter $\alpha>1$ constant in time. Up to the authors’ knowledge, scaled Poisson distributions have not yet been explored in the epidemiology literature, making this an original contribution. Altogether, the proposed weekly scaled Poisson epidemiological model writes